OpenStudy (anonymous):

A box with a square base and no top has a volume of 6 cubic meters. Express the total surface area, A, of the box as a function of the width, W, of the base?

OpenStudy (anonymous):

x=57.3 plusorminus 2 pi n

OpenStudy (anonymous):

sorry. i responded to the wrong person

OpenStudy (campbell_st):

so the square base has dimensions w and height h Area = w^2 then volume 64 = w^2h so \[h = \frac{64}{w^2}\] so in terms of surface area, the square base has and area \[A= w^2\] and the 4 rectangular sides \[A = t \times w \times h\] so the surface area is \[SA = w^2 + 4wh\] now subsitute h and you get \[SA = w^2 + \frac{64w}{w^2}\] or \[SA = w^2 + \frac{64}{w}\]

OpenStudy (campbell_st):

oops 4 rectangular sides have \[A = 4 \times w \times h\]

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

surface area without top means, \[SA = SA(base) + 4*SA(sides)\] assume width=x and height=y so volume of box= \[x^2y\] \[6/x^2=y\] Now substitute y value in above equation ....

OpenStudy (anonymous):

surface area of base is \[x^2\] and surface area of 4 sides =4*x*y

OpenStudy (anonymous):

Refer to the Mathematica solution attached.